The coefficient of linear expansion of steel is 11 × 10-6 per OC. A steel ball has a volume of

exactly 100 cm3 at 0OC. When heated to 100OC, what is its volume (in cm3)?


1
Expert's answer
2021-02-03T16:11:54-0500
"\dfrac{\Delta V}{V_0}=\alpha_V \Delta T,""V_1-V_0=3\alpha_LV_0 \Delta T,""V_1=V_0(1+3\alpha_L\Delta T),""V_1=100\ cm^3\cdot(1+3\cdot11\cdot10^{-6}\ ^{\circ}C^{-1}\cdot100\ ^{\circ}C)=100.33\ cm^3."